Consider a circuit containing five identical light
Answer #1

Bulb A = power = (I/2)^2 R = RI^2 / 4
Bulb B = power = (I/2)^2 R = RI^2 / 4
Bulb C = power = (I2/3)^2 R = 4RI^2 / 9
Bulb D = power = (I/3)^2 R = RI^2 / 9
Bulb E = power = (I/2)^2 R = RI^2 / 9

So
C > A = B > D = E